OpenStudy (anonymous):

help!?!? Find the length of the base of a square pyramid if the volume is 128 cubic inches and has a height of 6 inches. Use 3.14 for pi and round your answer to the nearest hundredth. 4 inches 8 inches 16 inches 24 inches

OpenStudy (anonymous):

Volume of a rectangular prism (or box) is: V = lwh {length x width x height} 8(x - 7)(x - 7) = 648 (x - 7)(x - 7) = 81 x² - 14x + 49 = 81 x² - 14x - 32 = 0 (x - 16)(x + 2) = 0 x - 16 = 0 or x + 2 = 0 x = 16 or x = -2 x = 16 x cannot be negative for x - 7 being the width and length the value of x is 16

hartnn (hartnn):

where did the prism come from :O

OpenStudy (raden):

Volume prism = 1/3 * base area * height

OpenStudy (anonymous):

oh ya....i just realized its not a prism.....thn idk what to do here

OpenStudy (raden):

opppss.. i mean pyramid lol

OpenStudy (raden):

1/3 * x^2 * 6 = 128 solve for x (x is the base of that pyramid)

OpenStudy (raden):

yeah, no pi here. because the base is a square not a circle

OpenStudy (anonymous):

how to find x by the way (finding the x was the reason i failed geometry )

OpenStudy (raden):

1/3 * x^2 * 6 = 128 see 1/3 * 6 = 2 so, we get 2 x^2 = 128 divided by 2 on both sides, get 2x^2/2 = 128/2 x^2 = 64 x = sqrt(64) = ... ?

OpenStudy (anonymous):

8

OpenStudy (raden):

yup, x = 8

OpenStudy (anonymous):

so the answer is 8 inches

OpenStudy (raden):

that's right!

OpenStudy (anonymous):

thx could u help me to solve another question??

OpenStudy (raden):

okay, close this and post it in the new question

OpenStudy (raden):

welcome :)